Actions
Ops/Comm Scheduling meeting¶
Location : CR11 R3 @ 15:30 pm on Tuesday 8th
Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402
Agenda¶
- Actions from last meeting
- Detector owners: Present ONE PLOT at this meeting that comments on the status of your detector.
- John G/Colin/Andy: LiH installation progress.
- Jaroslaw: Ramping/training plan for SSU/FC/SSD loosely agreed and evolving?
- DC: Reintroduce the dummy for searches
- Melissa/Ed: Define run settings with helium in one tracker and air in the other. Pass these on to Steve B., Chris R. and MOM. Update DC on tracker requirements for this run.
- Craig: Organise ISIS/MICE BPS tests in shutdown. Last carried out 18/3/15. Organising with Sheffield to their availability.
- Include Dean Adams & Mark Arnold.
- MOM report - Data-taking this week
- Lessons learnt during the run
- Summary of runs (MOM)
- Data plots from recent shift
- TOF (Durga) tof-online.pdf
- EMR (Francois) EMR_occupancy, EMR_beam_profile, EMR_range, EMR_PID
- Tracker (Melissa) Summary7696.pdf
- Cerenkov (?) CKOVA_vs_TOF01, CKOVB_vs_TOF01, CKOV_charge
- KL (?) KL_plots
- Review of web whiteboard : http://micewww.pp.rl.ac.uk/projects/operations/wiki/Web_Whiteboard
- DC report - activities in the Hall this week
- Focus coil and absorber installation status
- Controls and monitoring from last meeting
- Radiation monitor EPICS display freezes for ~2min
- Proton absorber requires manual resetting
- DAQ Heartbeat alarm
- Scheduling of required hall work and operations
- Installation of LiH
- Shifter training
- Safety talk and Hall walkaround : Tuesday morning
- PPE : Wednesday afternoon @ 3/4 pm (? - discuss)
- Beamline : Thursday afternoon @ 3/4 pm (? - discuss)
- AOB
- Chair of meetings for next week is DC/MOM
- Steve is teaching next week. DC will chair the meeting.
- DONM : 15/3/2016
Minutes
Present: Ryan, Ed, Henry, Melissa, Pierrick, Paul K, Paolo, Durga, Francois, Colin, Vicky B, Victoria B, Ken, John Nugent, Josef, Mark T, Craig, Jaroslaw, Chris R, Mike C, Scott, Alan Young, Yordan
Actions from last meeting
DC: LiH installation is in progress and on schedule for completion before Friday. Mylar windows in place
Jaroslaw: No further SSU or SSD training. Some possible training required for present hall FC (FC1) if used.
DC: Purchase dummy for hall search
Melissa/Ed: Define run settings with helium in one tracker and air in the other. Pass these on to Steve B., Chris R. and MOM. Presently He in both. Stands
Pierrick: Work with Craig work out if monitor or cable issue
Craig: Organising BPS test for machine physics run-up from 4th - 8th April. Need to do live ISIS trip
Actions arising from this meeting
DC: Bring web whiteboard up to date
Pierrick: Awaiting CDB settings for Proton absorber.
Paolo: Using Nagios for DAQ heartbeat alarm
Notes from this meeting
P.Kyberd reported a successful run period from last week. Smooth consistent operation of MICE.
Untrained personnel last week caused problems with controlled entry. Highlights importance of attending training and also justifies frequency of retraining.
DS cold box showing some abnormal instability so Mike and Mark plan to warm up and clean through with N2 and then cool again
Shifter training to be fitted in towards end of working week before weekend running
Yordan wants to test new prototype of online reconstruction. Requires 2hr debug window.
Mike C requires direct notice of desire to power existing FC in hall
Updated by MacWaters, Craig almost 8 years ago · 16 revisions